Please help me solve the problems of First-Degree Equations.
1) 7y - 7 = -1
7y-7+7 = -1 + 7
7y = 6
2) 5=4y/3+12
3) 2(x-3)= 6
4) 7x-4= 0
5) 3/2x+3 = 5
6) 3(x-5) = 15x
7) 4(x-1)-3(x+1) = 2
8) 4(x-2)- (3-x) = 3(x-1)
9) x/3 + x/4 + 7/12

Answer by roxy9229(32)

1) 7y - 7 = -1
7y-7+7 = -1 + 7
7y = 6
2) 5=4y/3+12
3) 2(x-3)= 6
4) 7x-4= 0
5) 3/2x+3 = 5
6) 3(x-5) = 15x
7) 4(x-1)-3(x+1) = 2
8) 4(x-2)- (3-x) = 3(x-1)
9) x/3 + x/4 + 7/12

1. 7y-7=-1
+7 +7
---------
7y= 6
- -
7 7
y= 6
-
7
2. not sure
3. 2(x-3)= 6
2x-6= 6
+6 +6
-----------
2x=12
- --
2 2
x=6
4. 7x-4= 0
+4 +4
----------
7x=4
- -
7 7
x=4/7
5-9...not sure